# Stabilizing and Solving Inverse Problems using Data and Machine Learning

### v2

@@ Abstract
We consider an inverse problem involving the reconstruction of the solution to a nonlinear partial differential equation (PDE) with unknown boundary conditions. Instead of direct boundary data, we are provided with a large dataset of boundary observations for typical solutions (collective data) and a bulk measurement of a specific realization. To leverage this collective data, we first compress the boundary data using proper orthogonal decomposition (POD) in a linear expansion. Next, we identify a possible nonlinear low-dimensional structure in the expansion coefficients using an autoencoder, which provides a parametrization of the dataset in a lower-dimensional latent space. We then train an operator network to map the expansion coefficients representing the boundary data to the finite element solution of the PDE. Finally, we connect the autoencoder’s decoder to the operator network which enables us to solve the inverse problem by optimizing a data-fitting term over the latent space. We analyze the underlying stabilized finite element method in the linear setting and establish an optimal error estimate in the $H^{1}$-norm. The nonlinear problem is then studied numerically, demonstrating the effectiveness of our approach.
@@ 1 Introduction
This is the situation that we wish to address in the present work. The objective is to construct a computational method that combines machine learning techniques for the data handling parts and hybrid network/finite element methods for the approximation in physical space. First, the large data set is mapped to a lower-dimensional manifold using an autoencoder or some other technique for finding low-dimensional structures, such as singular value decomposition or manifold learning. Then, we train a network to reproduce the solution map from the lower-dimensional set to the finite element space. Finally, this reduced order model solves a nonlinear inverse problem under the a priori assumption that the solution resides in a neighborhood of the lower-dimensional manifold.
@@
These developments are very useful in the context of inverse problems, where they have been utilized in both data- and model-driven inverse problems. In [32] a combination of networks and traditional methods is considered to recover the diffusion coefficient in Poisson’s and Burgers’ equations. In general, the same is done in [14] with the traditional method being FEM and the equations being elliptic and parabolic. Yet more examples of applying deep learning to this type of problem are given by [15, 4]. Anyone interested in the application of deep learning for PDE-solving has undoubtedly encountered PINNs [34] which are also used for inverse problems. Works not involving deep learning but still relevant are [28, 17] where projection-based reduced order models for inverse problems are presented. Taking the step to also include machine learning, some of the authors from the previous works give an overview of this mix in [24]. Another overview of using machine learning for inverse problems is given by [2]. In [23], an approach to reduce the error introduced by using operator learning for inverse problems is studied. As a contrast, [30] instead uses machine learning to reduce the error introduced by approximate forward models. Focusing instead on the other side of the computational spectrum, i.e., speed, [21] presents a physics-based deep learning methodology with applications to optimal control. The work [19] presents a modular machine learning framework for solving inverse problems in a latent space. Although using different techniques and approaches, this general description also holds for what we present here.
@@ 2.1 Inverse Problem
$\mathcal{G}=\Bigl{\{}g\in H^{s}(\partial\Omega)\,|\,g=\sum_{i=1}^{N}a_{i}\varphi_{i},\ a_{i}\in I_{i}\Bigr{\}}$ (2.3)
where $I_{i}$ are bounded intervals and $\varphi_{i}\in H^{s}(\partial\Omega)$. Below we will also consider access to a finite set $\mathcal{G}_{S}\subset\mathcal{G}$ of samples from $\mathcal{G}$,
@@ 2.2 Finite Element Method
for $10\leq m\leq k\leq p+$ 1. Here $\lesssim$ means that there is a positive constant in the inequality (typically on the right-hand side) and $N(T)$ is the union of all elements that share a node with $T$.
@@ 3.1 The Continuous Problem
This is a finite-dimensional problem, and therefore, existence follows from uniqueness. To prove uniqueness consider two solutions $u_{1}$ and $u_{2}$, we then have
@@ 3.2 The Discrete Problem
to represent linear combinations given coefficient vectors. By composing $I$ and $I_{h}$ with the coefficient extraction operator $\widehat{\cdot}$, we note that $v=I\widehat{v}$ for $v\in\mathcal{U}_{N}$ and $v=I_{h}\widehat{v}$ for $\mathcal{U}_{N,h}$. We also note that $I_{h}\widehat{v}$ is the Galerkin approximation defined by (3.9) of $v=I\widehat{v}$, since $\varphi_{n,h}$ is the Galerkin approximation of $\varphi_{n}$ for $n=1,\dots,N$, and we have the error estimate
@@
Now if we proceed as in (3.5) with the modes $\varphi_{n}$ replaced by the approximate modes $\varphi_{n,h}$, we can not directly use the same argument as in the continuous case to show that there is a unique solution since the discrete method does not possess the unique continuation property, and it doesn’t appear easy to quantify how small the mesh size must be to guarantee that the bound (3.8) holds on $\mathcal{U}_{N,h}$ as discussed in Remark 3.1.
Remark 3.1.
Note that the constant in (3.8) is characterized by the Rayleigh quotient
@@
$\boxed{m_{h}(v,w)=(v,w)_{\omega}+s_{\partial\Omega}(v,w)+s_{h}(v,w)}$ (3.23)
where
@@
controls the error at the boundary, and $s_{h}$ is the standard normal gradient jump penalty term
$s_{h}(v,w)=\sum_{F\in\mathcal{F}_{h}}h([\nabla v],[\nabla w])_{F}$ (3.25)
where $\mathcal{F}_{h}$ is the interior faces in the mesh $\mathcal{T}_{h}$. In the implementation, we estimate the $H^{1/2}(\partial\Omega)$-norm by
where $\nabla_{T}=(I-n\otimes n)\nabla$, with $n$ the unit normal to $\partial\Omega$ is the tangent derivative at $\partial\Omega$.
@@ 3.3 Error Estimates
    we obtain using Galerkin orthogonality followed by partial integration
    $((I-I_{h})\widehat{v},\psi)_{\Omega}$ $=a_{h}((I-I_{h})\widehat{v},\phi)$ (3.31)
    $=a_{h}((I-I_{h})\widehat{v},\phi-\pi_{h}\phi)$ (3.32)
@@
    where $\pi_{h}:H^{1}(\Omega)\rightarrow V_{h}$ is the interpolation operator and we used the standard trace inequality $\|w\|^{2}_{\partial T}\lesssim h^{-1}\|w\|_{T}^{2}+h\|\nabla w\|^{2}_{T}$ for $w\in H^{1}(T)$ on element $T\in\mathcal{T}_{h}$. Finally, using the elliptic regularity $\|\phi\|_{H^{2}(\Omega)}\lesssim\|\psi\|_{\Omega}$, we get
@@
    $\lesssim h(\|v\|_{s_{h}}+\|v\|_{h,\partial\Omega})$ (3.36)
@@
Define the stabilized projection,
$\boxed{u_{N,h}\in\mathcal{U}_{N,h}:\qquad(u_{N,h},v)_{m_{h}}=(u_{0},v)_{\omega}\qquad\forall v\in\mathcal{U}_{N,h}}$ (3.37)
@@
    $(u_{N},v)_{m_{h}}=(u_{N},v)_{\omega}=(u_{0},v)_{\omega}\qquad\forall v\in\mathcal{U}_{N}$ (3.42)
    We then have for any $v\in\mathcal{U}_{N,h}$,
    $(I_{h}\widehat{u}_{N}-u_{N,h},v)_{m_{h}}$ (3.43)
    $=(I_{h}\widehat{u}_{N},v)_{m_{h}}-(u_{0},v)_{\omega}$ (3.44)
    $=(I_{h}\widehat{u}_{N}-u_{N},v)_{m_{h}}+(u_{N},v)_{m_{h}}-(u_{0},v)_{\omega}$ (3.45)
    $=(I_{h}\widehat{u}_{N}-u_{N},v)_{m_{h}}+(u_{N},v)_{\omega}-(u_{0},v)_{\omega}$ (3.46)
    $=(I_{h}\widehat{u}_{N}-u_{N},v)_{m_{h}}+(u_{N},v-I\widehat{v})_{\omega}-(u_{0},v-I\widehat{v})_{\omega}$ (3.47)
@@
    where we used (3.42) in (3.42); the definition (3.5) of $u_{N}$ to subtract $I\widehat{v}\in\mathcal{U}_{N}$ in (3.46); the stability (3.27); the bounds $\|\widehat{u}_{N}\|_{\mathbb{R}^{N}}\lesssim\|u_{N}\|_{\omega}$ and $\|u_{N}\|_{\omega}\lesssim\|u_{0}\|_{\omega}$. Thus, we conclude that
@@
where $(r,\theta)$ are the standard polar coordinates. Let $\omega$ be the disc centered at the origin with radius $r_{\omega}$. We note that when $n$ becomes large, the modes become small in the disc $\omega$, and therefore, the inverse problem becomes increasingly ill-posed, for instance, the constant in an estimate of the type
@@
and thus becomes arbitrarily large when $n$ becomes large. But, if we, from observations, can conclude that only modes with $n<n_{g}$ for some $n_{g}$ are present, then the stability is controlled. Note also that the stability is directly related to where the disc $\omega$ is placed. If it is located close to the boundary, the stability improves.

### v3

# Stabilizing and Solving Unique Continuation Problems by Parameterizing Data and Learning Finite Element Solution Operators

@@ Abstract
We consider an inverse problem involving the reconstruction of the solution to a nonlinear partial differential equation (PDE) with unknown boundary conditions. Instead of direct boundary data, we are provided with a large dataset of boundary observations for typical solutions (collective data) and a bulk measurement of a specific realization. To leverage this collective data, we first compress the boundary data using proper orthogonal decomposition (POD) in a linear expansion. Next, we identify a possible nonlinear low-dimensional structure in the expansion coefficients using an autoencoder, which provides a parametrization of the dataset in a lower-dimensional latent space. We then train an operator network to map the expansion coefficients representing the boundary data to the finite element (FE) solution of the PDE. Finally, we connect the autoencoder’s decoder to the operator network which enables us to solve the inverse problem by optimizing a data-fitting term over the latent space. We analyze the underlying stabilized finite element method (FEM) in the linear setting and establish an optimal error estimate in the $H^{1}$-norm. The nonlinear problem is then studied numerically, demonstrating the effectiveness of our approach.
@@ 1 Introduction
This is the situation that we wish to address in the present work. The objective is to construct a computational method that combines machine learning techniques for the data handling parts and hybrid network/finite element methods (FEMs) for the approximation in physical space. First, the large data set is mapped to a lower-dimensional manifold using an autoencoder or some other technique for finding low-dimensional structures, such as singular value decomposition or manifold learning. Then, we train a network to reproduce the solution map from the lower-dimensional set to the finite element space. Finally, this reduced order model solves a nonlinear inverse problem under the a priori assumption that the solution resides in a neighborhood of the lower-dimensional manifold.
@@
These developments are very useful in the context of inverse problems, where they have been utilized in both data- and model-driven inverse problems. In [37] a combination of networks and traditional methods is considered to recover the diffusion coefficient in Poisson’s and Burgers’ equations. In general, the same is done in [14] with the traditional method being FEM and the equations being elliptic and parabolic. Yet more examples of applying deep learning to this type of problem are given by [15, 4]. Anyone interested in the application of deep learning for PDE-solving has undoubtedly encountered Physics-Informed Neural Networks (PINNs) [39] which are also used for inverse problems. Works not involving deep learning but still relevant are [31, 17] where projection-based reduced order models for inverse problems are presented. Taking the step to also include machine learning, some of the authors from the previous works give an overview of this mix in [26]. Another overview of using machine learning for inverse problems is given by [2]. In [25], an approach to reduce the error introduced by using operator learning for inverse problems is studied. As a contrast, [33] instead uses machine learning to reduce the error introduced by approximate forward models. Focusing instead on the other side of the computational spectrum, i.e., speed, [21] presents a physics-based deep learning methodology with applications to optimal control. The work [19] presents a modular machine learning framework for solving inverse problems in a latent space. Although using different techniques and approaches, this general description also holds for what we present here.
Comparison between this Work and Others.
To the best of our knowledge, there are no previous works in the literature with a similar theoretical foundation addressing this type of data assimilation problem. Notably, the importance of the finite dimensionality of boundary data for stability, and thus the necessity to reduce the dimension of measured population data as much as possible, has only been considered in [5] using classical methods. Here, we apply autoencoders and operator learning to this problem for the first time. To provide context on how other approaches might perform compared to ours, we note that the stabilized method proposed here yields optimally converging approximations, contingent on the properties of the finite element (FE) space and the stability of the inverse problem. This is not the case for Tikhonov regularized approaches, where discretization is typically applied without further consideration of numerical stability. Bayesian inference methods usually share a similar shortcoming, depending on the choice of prior. We note that in our computational examples, stabilization was not necessary, indicating that the space discretization was sufficiently well-resolved. The use of PINNs in this context leads to a formulation where the strong form of the PDE is minimized. This presents complications, as boundary conditions are generally difficult to impose in network approximations, particularly on the finite-dimensional subspace. Additionally, there appears to be no way to eliminate spurious local minima in the PDE approximation when using PINNs. In our case, since we minimize a convex functional over the finite element space for all parameter values, the space discretization part does not suffer from this defect. Nonetheless, the optimization could converge to local minima when networks approximate the operator, a common shortcoming with network approximation methods.
Concerning the approach to learning the physical model, the method we use is presented in detail in [30], where the focus lies on the method itself as opposed to here, where the focus is on applying it to inverse problems. In [30], a comparison with other machine learning approaches is made so we refer to this work for details and only give a brief characterization here:
  • The core concept is to learn a finite element solution operator. The output is thus an approximation of a finite element solution. An advantage of this is that the method can be combined with standard FEM for support and enhancement in both theory and practice.
  • A multilayer perceptron (MLP) is used to approximate the solution operator.
  • The finite element part enters by using a mesh and basing the loss function on an energy functional that when minimized gives the FE-solution. An alternative is to use the weak residual, which although is more general seems to be more computationally costly.
  • The input to the network is a parameterization of problem data, e.g., right-hand side functions and boundary values. The network thus learns a parameterized family of PDE-problems as opposed to only a single problem.
  • The method is by default data-free, meaning no input-output data pairs. Instead input is sampled from probability distributions. However, the method allows for the incorporation of data sets as demonstrated here.
None of these individual features is new in physics-based deep learning, but to the best of our knowledge, this specific combination of them has not been studied outside of this work and [30].
Looking through the literature for other works employing deep learning for unique continuation problems, we find two different types. The first type presented in [27] is a data-driven approach for parameterizing both boundary conditions and solutions for flow problems in a cylinder. The velocity distribution is observed in a downstream cross section and the objective is to find a matching inlet profile. Although numerical PDE simulations are used to generate data, this learning approach is physics-free in the contextual sense. The second more common type uses PINNs and seeks the full solution given pointwise observations in a subdomain of the solution domain. In [35] four standard linear problems (Poisson’s equation, the heat and wave equations, and Stokes flow) are considered. For a nonlinear problem, see [22] where the 2D Navier–Stokes equations are studied. The work [36] again considers a linear problem, the Helmholtz equation. A drawback of these PINNs-based works is that the neural networks only learn a single PDE solution during training. Comparing these works with ours, we see some differences with our approach: First, it is physics-based in contrast to [27]. Second, it allows for learning an entire class of related PDE solutions as opposed to a single one as in the PINNs-based works. With these points in mind, we think that our deep learning approach to unique continuation problems provides a novelty that can further the field.
@@
Notation.
  • We use $\lesssim$ to mean that there is a positive constant in the inequality (typically on the right-hand side).
  • For a bounded domain or a set of mesh features $D$, we denote by $\|\cdot\|_{D}$ and $(\cdot,\cdot)_{D}$ the standard $L^{2}(D)$-norm and inner product, respectively. Some common instances in the text are $D=\omega,\Omega,\partial\Omega,\mathcal{F}_{h}$.
  • We denote by $\|\cdot\|_{\mathbb{R}^{N}}$ the standard absolute value for vectors in $\mathbb{R}^{N}$. We note that it should not be confused with the $L^{2}$-notation in the previous point. The reason for using the notation $\|\cdot\|_{\mathbb{R}^{N}}$ is because we use it on the expansion coefficients of functions, thus making expressions involving norms on both functions and their coefficients more general and consistent.
  • For a positive-definite bilinear form $B$ we denote the corresponding norm by $\|\cdot\|_{B}$, i.e., $\|v\|_{B}^{2}:=B(v,v)$. An example from the text is $B=m_{h}$.
@@ 2.1 Inverse Problem
$\mathcal{G}=\Bigl{\{}g\in H^{1/2}(\partial\Omega)\,|\,g=\sum_{i=1}^{N}a_{i}\varphi_{i},\ a_{i}\in I_{i}\Bigr{\}}$ (2.3)
where $I_{i}$ are bounded intervals and $\varphi_{i}\in H^{1/2}(\partial\Omega)$. Below we will also consider access to a finite set $\mathcal{G}_{S}\subset\mathcal{G}$ of samples from $\mathcal{G}$,
@@ 2.2 Finite Element Method
for $10\leq m\leq k\leq p+$ 1. Here $N(T)$ is the union of all elements that share a node with $T$.
@@ 3.1 The Continuous Problem
This is a finite-dimensional problem, and therefore, existence follows from uniqueness. To prove uniqueness, consider two solutions $u_{1}$ and $u_{2}$, we then have
@@ 3.2 The Discrete Problem
to represent linear combinations given coefficient vectors, where $\mathcal{U}_{N,h}=\text{span}\{\varphi_{n,h}\}_{n=1}^{N}$. By composing $I$ and $I_{h}$ with the coefficient extraction operator $\widehat{\cdot}$, we note that $v=I\widehat{v}$ for $v\in\mathcal{U}_{N}$ and $v=I_{h}\widehat{v}$ for $\mathcal{U}_{N,h}$. We also note that $I_{h}\widehat{v}$ is the Galerkin approximation defined by (3.9) of $v=I\widehat{v}$, since $\varphi_{n,h}$ is the Galerkin approximation of $\varphi_{n}$ for $n=1,\dots,N$, and we have the error estimate
@@
Now if we proceed as in (3.5) with the modes $\varphi_{n}$ replaced by the approximate modes $\varphi_{n,h}$, we can not directly use the same argument as in the continuous case to show that there is a unique solution since the discrete method does not possess the unique continuation property, and it doesn’t appear easy to quantify how small the mesh size must be to guarantee that the bound (3.8) holds on $\mathcal{U}_{N,h}$.
To quantify the discrete stability, note that the constant in (3.8) is characterized by the Rayleigh quotient
@@
$\boxed{m_{h}(v,w)=(v,w)_{\omega}+s_{h,\partial\Omega}(v,w)+s_{h}(v,w)}$ (3.23)
Here
where $\nabla_{T}=(I-n\otimes n)\nabla$ is the tangential derivative on $\partial\Omega$ with $n$ denoting the unit normal to $\partial\Omega$. The form $s_{h}$ is the standard normal gradient jump penalty term
$s_{h}(v,w)=\sum_{F\in\mathcal{F}_{h}}h([\nabla v],[\nabla w])_{F}$ (3.25)
where $\mathcal{F}_{h}$ is the interior faces in the mesh $\mathcal{T}_{h}$. The role of the form $s_{h,\partial\Omega}$ is to give control of the distance of the approximation to the finite dimensional set $\mathcal{G}$ in the $H^{1/2}(\partial\Omega)$-norm. In principle the form
@@
could be used directly, but to obtain a stabilization term that is easier to handle in practice we note that by the Galigliardo-Nirenberg inequality, $\|v\|_{H^{1/2}(\partial\Omega)}\lesssim\|v\|_{L^{2}(\partial\Omega)}^{1/2}\|v\|_{H^{1}(\partial\Omega)}^{1/2}$, there holds $s_{\partial\Omega}(v,v)\lesssim s_{h,\partial\Omega}(v,v)$, which is sufficient for stability.
@@ 3.3 Error Estimates
    we obtain by partial integration followed by Galerkin orthogonality
    $((I-I_{h})\widehat{v},\psi)_{\Omega}=((I-I_{h})\widehat{v},\mathcal{P}\phi)_{\Omega}=a_{h}((I-I_{h})\widehat{v},\phi)=a_{h}((I-I_{h})\widehat{v},\phi-\pi_{h}\phi)$ (3.31)
    where $\pi_{h}:H^{1}(\Omega)\rightarrow V_{h}$ is the interpolation operator. Performing another partial integration, we get
    $a_{h}((I-I_{h})\widehat{v},\phi-\pi_{h}\phi)$ (3.32)
@@
    where we used the standard trace inequality $\|w\|^{2}_{\partial T}\lesssim h^{-1}\|w\|_{T}^{2}+h\|\nabla w\|^{2}_{T}$ for $w\in H^{1}(T)$ on an element $T\in\mathcal{T}_{h}$. Finally, using the elliptic regularity $\|\phi\|_{H^{2}(\Omega)}\lesssim\|\psi\|_{\Omega}$, combining the results, and taking $\psi=(I-I_{h})\widehat{v}$, we get
@@
    $\lesssim h(\|v\|_{s_{h}}+\|v\|_{s_{h,\partial\Omega}})$ (3.36)
@@
We define the stabilized projection,
$\boxed{u_{N,h}\in\mathcal{U}_{N,h}:\qquad m_{h}(u_{N,h},v)=(u_{0},v)_{\omega}\qquad\forall v\in\mathcal{U}_{N,h}}$ (3.37)
@@
    $m_{h}(u_{N},v)=(u_{N},v)_{\omega}\qquad\forall v\in\mathcal{U}_{N,h}$ (3.42)
    Then by subtracting and adding $u_{N}$ in the first argument to $m_{h}$, we have for any $v\in\mathcal{U}_{N,h}$,
    $m_{h}(I_{h}\widehat{u}_{N}-u_{N,h},v)$ (3.43)
    $=m_{h}(I_{h}\widehat{u}_{N}-u_{N},v)+m_{h}(u_{N},v)-m_{h}(u_{N,h},v)$ (3.44)
    $=m_{h}(I_{h}\widehat{u}_{N}-u_{N},v)+(u_{N},v)_{\omega}-(u_{0},v)_{\omega}$ (3.45)
    $=m_{h}(I_{h}\widehat{u}_{N}-u_{N},v)+(u_{N},v-I\widehat{v})_{\omega}-(u_{0},v-I\widehat{v})_{\omega}$ (3.46)
    where we used (3.42) and (3.37) on the second and third terms in (3.44), respectivley, and the definition (3.5) of $u_{N}$ to subtract $I\widehat{v}\in\mathcal{U}_{N}$ in (3.45). Employing continuity of the involved forms, we get
    $m_{h}(I_{h}\widehat{u}_{N}-u_{N,h},v)$ (3.47)
@@
    where we used the stability (3.27) and the bounds $\|\widehat{u}_{N}\|_{\mathbb{R}^{N}}\lesssim\|u_{N}\|_{\omega}$ and $\|u_{N}\|_{\omega}\lesssim\|u_{0}\|_{\omega}$. Thus by taking $v=I_{h}u_{N}-u_{N,h}$, we conclude that
@@
where $(r,\theta)$ are the standard polar coordinates. Let $\omega$ be the disc centered at the origin with radius $r_{\omega}$. We note that when $n$ becomes large, the modes become small in the disc $\omega$, and therefore, the inverse problem becomes increasingly ill-posed, see Figure 2(a). For instance, the constant in an estimate of the type
@@
and thus becomes arbitrarily large when $n$ becomes large. But, if we, from observations, can conclude that only modes with $n<n_{g}$ for some $n_{g}$ are present, then the stability is controlled, see Figure 2(b). Note also that the stability is directly related to where the disc $\omega$ is placed. If it is located close to the boundary, the stability improves.
(a) Domain and first five modes of $\varphi_{2n+1}=r^{n}\sin(n\theta)$
(b) Scaling of stability constants
Figure 2: Illustrations for the analytical example with exponential growth of the stability constant. In (a), we show the unit disc domain containing a subdomain $\omega$, in the form of a centered disc of radius $r_{\omega}$. Looking at the first five non-zero modes $\varphi_{2n+1}=r^{n}\sin(n\theta)$ from the expansion (3.60) we see that these modes rapidly becomes very small within $\omega$, making the problem of retrieving the coefficient values in the expansion based on the solution within $\omega$ increasingly ill-posed for larger $n$. In (b), we illustrate how the constants in the stability estimate (3.61) scales exponentially with $n$ for different values of the radius $r_{\omega}$. By utilizing observations, we may conclude an upper bound on $n$, which in turn puts an upper bound on the size of the stability constant.
@@ 4.1 Processing the Boundary Data
We combine linear and nonlinear dimensionality reduction techniques by first using PCA on the data to get a POD-basis and then using autoencoders on the POD-coefficients for further reduction. Such combinations are not uncommon, see, e.g., [24], and there are several reasons why we do this: In general, an initial linear reduction may function as a relatively cheap preprocessing step to aid a subsequent nonlinear reduction that typically is more expensive. More specifically, we consider methodology for progressing from a fully linear problem (linear PDE, linear data) to a fully nonlinear one (nonlinear PDE, nonlinear data), hence both linear and nonlinear techniques that can be combined are required. The reason for using POD is that it very cheaply and naturally gives a basis that can be used for both the linear and nonlinear PDE-solving techniques considered here. The reason for using autoencoders is simply because they have the same general network architecture already used for the operator networks which also makes combinations with them seem natural.